OpenStudy (anonymous):

Which of the following is a prime factorization of 84? A. 2 x 2 x 3 x 7 B. 4 x 3 x 7 C. 2 x 2 x 21 D. 2 x 2 x 2 x 7

OpenStudy (jadedry):

Prime numbers are numbers that can only be divided by 1 and it'self. The exception to this rule is the number 1. 84 = 2 * 42 = 2 * 2 * 21 = 2 * 2 * 3 * 7 Therefore the answer is A.

OpenStudy (anonymous):

Thank You!